# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v2

@@ IV-B2 One-stage 3D Visual Grounding methods
3D-SPS [142] takes the 3D VG task as a keypoint selection problem and avoids the separation of detection and matching. Specifically, 3D-SPS initially coarsely samples language-related keypoints through the description-aware keypoint sampling module. Subsequently, it finely selects target keypoints and predicts the foundation using the goal-oriented progressive mining module. Inspired by the 2D image language pre-train model such as MDETR[156] and GLIP[157], BUTD-DETR[143] proposes a bottom-up top-down detection transformer that can be used for 2D and 3D VG. Concretely, BUTD-DETR utilizes labeled bottom-up box proposals and top-down language descriptions to guide the decoding of target objects and corresponding language spans through the prediction head.
However, the aforementioned methods either extract sentence-level features that couple all words or focus more on object names in the description, which would lose the word-level information or neglect other attributes. To address these issues, EDA[144] explicitly decouples the textual attributes in a sentence and conducts dense alignment between such fine-grained language and point cloud objects. Firstly, the long text is decoupled into five semantic components, including main object, auxiliary object, attributes, pronoun, and relationship. Subsequently, the dense alignment is designed to align all object-related decoupled textual semantic components with visual features. To be able to reason human intentions from implicit instructions, ReGround3D [145] designs a visual-centric reasoning module, powered by a Multi-modal Large Language Model, and a 3D grounding module that accurately obtains object locations by revisiting enhanced geometry and fine-grained details from 3D scenes. Additionally, a Chain-of-Grounding mechanism is employed to improve 3D reasoning grounding through interleaved reasoning and grounding steps.
@@ IV-C Visual Language Navigation
Below, we will introduce visual-linguistic navigation divided into two parts: datasets and methods.
@@ IV-C1 Datasets
In visual-linguistic navigation, natural language instructions can be a series of detailed action descriptions, a fully described goal, or just a roughly described task, even only the demands of human. The tasks that embodied agents need to complete maybe just a single navigation, or navigation with interaction, or multiple navigation tasks that need to be completed in sequence. These differences bring different challenges to visual-linguistic navigation, and many different datasets have been born as a result. Based on these differences, we will introduce the important datasets in visual-linguistic navigation separately.
Room to Room [65] is a visual-linguistic navigation dataset created by Anderson et al. based on Matterport3D. In R2R, embodied agents navigate according to step-by-step instructions, choosing the next adjacent navigation graph node to advance based on visual observations until they reach the target location. Embodied agents need to dynamically track progress to align the navigation process with fine-grained instructions. Room-for-Room [158] extends the paths in R2R to longer trajectories, requires higher for the long-distance instruction and history alignment capabilities of embodied agents. VLN-CE [159] extends R2R and R4R to continuous environments, embodied agents can move freely in the scene. This makes the action decision of embodied agents more difficult. Different from the above datasets based on indoor scenes, Chen et al. created the TOUCHDOWN dataset [160] based on Google Street View. In TOUCHDOWN, embodied agents follow instructions to navigate in the street view rendering simulation of New York City to find the specified object.
@@
Shridhar et al. proposed the ALFRED dataset [164] based on the AI2-THOR simulator. In ALFRED, embodied agents need to understand environmental observations and complete household tasks in an interactive environment according to coarse-grained and fine-grained instructions. The core task of OVMM [165] is to pick any object in any unseen environment and place it in a specified location. Embodied agents need to locate the target object in the home environment, navigate and grab it, and then navigate to the target location to put down the object. OVMM provides a simulation based on Habitat and a framework for implementation in the real world. Behavior-1K [166], based on the survey of human needs, designed 1000 long-sequence, complex skill-dependent daily tasks in OmniGibson, which is an extension of iGibson. Embodied agents need to complete long-span navigation-interaction tasks which may contain thousands of low-level action steps based on visual information and natural language instructions. Such complex tasks pose higher requirements for the understanding and memory of embodied agents.
@@ IV-C2 Method
Different learning schemes also explore how to utilize historical trajectories and memories better. Through adversarial learning, CMG [170] alternates between imitation learning and exploration encouragement schemes, effectively strengthening the understanding of instructions and historical trajectories, shortening the difference between training and inference. GOAT [178] directly trains unbiased models through Backdoor Adjustment Causal Learning (BACL) and Frontdoor Adjustment Causal Learning (FACL), conducts contrastive learning with vision, navigation history, and their combination to instructions, enabling the agent to make fuller use of information. The enhanced cross-modal matching method proposed by RCM [171] uses goal-oriented external rewards and instruction-oriented internal rewards to perform cross-modal grounding globally and locally and learns from its own historical good decisions through self-supervised imitation learning. FSTT [176] introduces TTA into visual-linguistic navigation and optimizes the model in terms of gradients and model parameters at two scales of time steps and tasks, effectively improving model performance.
The specific application of large models in Memory-Understanding based methods is to understand the representation of historical memory and to understand the environment and tasks based on its extensive world knowledge. NaviLLM [175] integrates the historical observation sequence into the embedding space through the visual encoder, inputs the multi-modal information of the fusion encoding into the large language model and fine-tunes it, reaching the state-of-the-art on multiple benchmarks. NaVid [180] makes improvements in the encoding of historical information, achieves different degrees of information retention on historical observations and current observations through different degrees of pooling. DiscussNav [177] assigns large model experts with different abilities to different roles, drives the large models to discuss before navigation actions to complete navigation decisions, and achieves excellent performance in zero-shot visual-linguistic navigation.
@@
Some reinforcement learning methods are also applied to predict and explore future states. LookBY [181] employs reinforcement prediction to enable the prediction module to imitate the world and forecast future states and rewards. This allows the agent to directly map “current observations” and “predictions of future observations” to actions, achieving state-of-the-art performance at the time. The rich world knowledge and zero-shot performance of large models provide many possibilities for Future-Prediction based methods. MiC [184] requires the large language model to directly predict the target and its possible location from the instructions and provides navigation instructions through the description of scene perception. This method requires the large language model to fully exert its ‘imagination’ and build an imagined scene through prompts.
@@
IV-D Non-Visual Perception: Tactile
The skin facilitates human tactile perception. The skin changes shape when touched, and its abundant nerve cells send electrical signals [189]. This tactile perception allows humans to grasp handy work fully. Therefore, touch is vital for robots to interact with the real world. A sense of touch enables robots to acquire information such as material, shape, temperature, and even objects’ contact force and gravity. Current work on tactile perception focuses on three areas: sensor design, dataset construction, and application. Tactile perception undoubtedly enhances the human-computer interaction experience and holds great promise[190, 191, 192].
@@ IV-D1 Sensor Design
Tactile sensor design methods can be divided into three categories: non-vision-based, vision-based, and multi-modal. At early time, tactile sensors were chiefly engineered to register fundamental, low-dimensional sensory outputs such as force, pressure, vibration, and temperature[193, 194, 195, 196, 197, 198]. Their principles are mostly related to electricity and physical mechanics and their data is often low-dimension series with temporal correlations. One of the notable representatives is BioTac [199] and its simulator[200]. Since computer vision had a wonderful performance, there has been a growing focus on vision-based tactile sensors which obtains tactile through optical principles. Using images of the gel’s deformation as tactile information, vision-based tactile sensors such as GelSight[201], Gelslim[202], DIGIT[203], 9DTact[204], TacTip[205], GelTip[206] and AllSight[207] have been used for numerous applications. At the same time, the simulation of tactile sensors like TACTO[208] and Taxim[209], incorporating elastomer deformation together with the optical simulation, are also widely used. Recent work has focused on low costs[204] and the installation on robotic hands[203, 210, 211]. Recently, multi-modal tactile sensors are also emerging. Inspired by the touch mechanism of human skin, works progressed in multi-modal tactile skin which comprised multi information like pressure, proximity, acceleration and temperature detectors. Solutions often involve techniques like flexible materials and modular design.
@@ IV-D2 Datasets
The datasets of non-vision sensors are contained with electrode values 3D net force vectors, and contact location. Therefore, the objects in the datasets are usually force samples and grasping samples. Its tasks are mainly the estimation of force types, force value and grasping details. The datasets mainly collected by BioTac series[199]. As for vision-based sensors, with their high-resolution images of deformation gel, except estimating the force information and sliding, they focus more on texture classification and 3D reconstruction. The objects in the datasets are usually household object, wildlife environments, different materials and grasping items.At the same time, as the image information can be easily aligned and bound with the other modalities (images, language, audio, etc)[16, 212], the perception of touch in embodied agent mainly revolves around visual-based sensors. The datasets revolves around the Geisight sensors, DIGIT sensor and their simulators[201, 203, 208, 204]. We introduce ten commonly used tactile datasets, which are summarized in Table VIII.
@@
TABLE VIII: Comparison of different vision-based tactile datasets.
@@ IV-D3 Methods
Tactile perception has numerous applications. The information acquired from sensor enables robots to perform accurate robotic manipulation task, finish multi-modal work and even enhance their abilities in 3D reconstruction and localization.
1. Robotic Manipulation. In these task, bridging the sim-to-real gap is more than important. Reinforcement learning and GAN-based methods have been proposed to address variations in accurate, on-time robotic manipulation tasks.
@@
GAN-based method. ACTNet[225] proposed an unsupervised adversarial domain adaptation method to narrow the domain gap for pixel-level tactile perception tasks. An adaptively correlative attention mechanism was introduced to improve the generator, which is capable of leveraging global information and focusing on salient regions. However, pixel-level domain adaptation lead to error accumulation, degrade performance, and increased structural complexity and training costs. In comparison, STR-Net[226] proposed a feature-level unsupervised framework for tactile images, narrowing the domain gap for feature-level tactile perception tasks.
Moreover, some methods focuse on sim-to-real. For example, the Tactile Gym 2.0 [227]. However, due to its complexity and high cost, it is challenging for practical application.
2. Classification&Recognition. Tactile representations learning focused on material classification work and multi-modal understanding. Method can be divied into 2 categories:Traditional Methods and LLMs&VLMs Methods.
@@
3. 3D Reconstruction. Suresh et al. [238] incrementally reconstructed the local shape of 3D household objects from a sequence of tactile images and a noisy depth map, representing the 3D shape as a signed distance function sampled from a Gaussian process and reformulated it as probabilistic inference on a spatial graph, providing a robust method for local shape reconstruction. Smith et al. [239] presented an effective chart-based approach. It took grasping parameters, a group of touch readings and an RGB image of the object as input and encoded them to charts, using a neural network to reconstruct 3D shape estimate. Then, the grop updated them into active touch learned models with sliding sensors[240]. Comi et al. [241] employed deep learning for 3D shape reconstruction solely based on tactile input. They used a CNN to map tactile images into local meshes and a DeepSDF-based model to predict the complete 3D shape.
@@ IV-D4 Difficulties
1) Disadvantages of sensors with different principles: traditional sensors provides simple, limited and low-dimension data, posing challenges for multi-modal learning. Meanwhile, vision-based sensors and electronic skins, although offering high accuracy, are cost-prohibitive. And vision-based sensors are unable to provide temperature information. 2) Difficulties in data acquisition: tactile datasets are rare and heterogeneous, lacking the standardized, extensive repositories found in fields like vision. Additionally, data collection is difficult. It is difficult to gather both tactile and visual information together, although some efforts have been made in simplified collection devices. 3) Difficulties in inconsistent standards: there are a variety of sensors on the market, with inconsistent standards and principles. Even with similar imaging patterns, vision-based tactile sensors’ design and calibration still results in a significant domain gap. Unlike standardized formats for visual and auditory data, tactile data formats vary widely among sensor manufacturers, which makes it difficult for large-scale learning on data collected from heterogeneous sensors, limiting the usefulness of publicly available tactile datasets.
@@ V-A Embodied Question Answering
For Embodied Question Answering (EQA) task, the agent needs to explore the environment from a first-person perspective to gather information necessary to answer the given questions. An agent with autonomous exploration and decision-making capabilities must not only consider which actions to take to explore the environment but also determine when to stop exploring to answer questions. Existing works focus on different types of questions, some of which are shown in Fig. 11. In this section, we first introduce the existing datasets and then discuss related methods.
@@ V-A2 Methods
The EQA task was first proposed by Das et al.[242]. In their work, the agent consists of four main modules: vision, language, navigation, and answering. These modules are primarily constructed using traditional neural building blocks - Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs). These modules undergo training in two phases. Initially, the navigation and response modules are trained independently on automatically generated expert navigation demonstrations using imitation or supervised learning. Subsequently, in the second phase, the navigation architecture is fine-tuned using policy gradients. Some subsequent works[256, 257] retain modules like the question answering module proposed by Das et al.[242] and improve the model. Additionally, Wu et al.[257] proposed integrating the navigation and QA modules into a unified SGD training pipeline for joint training, thereby avoiding employing deep reinforcement learning to simultaneously train the separately trained navigation and question answering modules.

### v3

@@ IV-B2 One-stage 3D Visual Grounding methods
3D-SPS [146] took the 3D VG task as a keypoint selection problem and avoided the separation of detection and matching. Specifically, 3D-SPS initially coarsely samples language-related keypoints through the description-aware keypoint sampling module. Subsequently, it finely selects target keypoints and predicts the foundation using the goal-oriented progressive mining module. Inspired by the 2D image language pre-train model such as MDETR[160] and GLIP[161], BUTD-DETR[147] proposed a bottom-up top-down detection transformer that can be used for 2D and 3D VG. Concretely, BUTD-DETR utilizes labeled bottom-up box proposals and top-down language descriptions to guide the decoding of target objects and corresponding language spans through the prediction head.
However, the aforementioned methods either extract sentence-level features that couple all words or focus more on object names in the description, which would lose the word-level information or neglect other attributes. To address these issues, EDA[148] explicitly decoupled the textual attributes in a sentence and conducted dense alignment between such fine-grained language and point cloud objects. Firstly, the long text is decoupled into five semantic components, including main object, auxiliary object, attributes, pronoun, and relationship. Subsequently, the dense alignment is designed to align all object-related decoupled textual semantic components with visual features. To be able to reason human intentions from implicit instructions, ReGround3D [149] designed a visual-centric reasoning module, powered by a Multi-modal Large Language Model, and a 3D grounding module that accurately obtains object locations by revisiting enhanced geometry and fine-grained details from 3D scenes. Additionally, a Chain-of-Grounding mechanism is employed to improve 3D reasoning grounding through interleaved reasoning and grounding steps.
@@ IV-C Visual Language Navigation
Below, we will introduce VLN divided into two parts: datasets and methods.
@@ IV-C1 Datasets
In VLN, natural language instructions can be a series of detailed action descriptions, a fully described goal, or just a roughly described task, even only the demands of human. The tasks that embodied agents need to complete maybe just a single navigation, or navigation with interaction, or multiple navigation tasks that need to be completed in sequence. These differences bring different challenges to VLN, and many different datasets have been built. Based on these differences, we introduce some important VLN datasets.
Room to Room [65] is a VLN dataset created by Anderson et al. based on Matterport3D. In R2R, embodied agents navigate according to step-by-step instructions, choosing the next adjacent navigation graph node to advance based on visual observations until they reach the target location. Embodied agents need to dynamically track progress to align the navigation process with fine-grained instructions. Room-for-Room [162] extends the paths in R2R to longer trajectories, requires higher for the long-distance instruction and history alignment capabilities of embodied agents. VLN-CE [163] extends R2R and R4R to continuous environments, embodied agents can move freely in the scene. This makes the action decision of embodied agents more difficult. Different from the above datasets based on indoor scenes, Chen et al. created the TOUCHDOWN dataset [164] based on Google Street View. In TOUCHDOWN, embodied agents follow instructions to navigate in the street view rendering simulation of New York City to find the specified object.
@@
Shridhar et al. [168] proposed the ALFRED dataset based on the AI2-THOR simulator. In ALFRED, embodied agents need to understand environmental observations and complete household tasks in an interactive environment according to coarse-grained and fine-grained instructions. The task of OVMM [169] is to pick any object in any unseen environment and place it in a specified location. Embodied agents need to locate the target object in the home environment, navigate and grab it, and then navigate to the target location to put down the object. OVMM provides a simulation based on Habitat and a framework for implementation in the real world. Behavior-1K [170] is a dataset based on human needs, comprising 1,000 long-sequence, complex, skill-dependent daily tasks, which are designed within OmniGibson, an extension of the iGibson simulation environment. Embodied agents need to complete long-span navigation-interaction tasks which may contain thousands of low-level action steps based on visual information and natural language instructions. Such complex tasks pose higher requirements for the understanding and memory of embodied agents.
@@ IV-C2 Method
Different learning schemes explore how to utilize historical trajectories and memories better. Through adversarial learning, CMG [174] alternates between imitation learning and exploration encouragement schemes, effectively strengthening the understanding of instructions and historical trajectories, shortening the difference between training and inference. GOAT [182] directly trains unbiased models through Backdoor Adjustment Causal Learning (BACL) and Frontdoor Adjustment Causal Learning (FACL), conducts contrastive learning with vision, navigation history, and their combination to instructions, enabling the agent to make fuller use of information. The enhanced cross-modal matching method proposed by RCM [175] uses goal-oriented external rewards and instruction-oriented internal rewards to perform cross-modal grounding globally and locally and learns from its own historical good decisions through self-supervised imitation learning. FSTT [180] introduces TTA into VLN and optimizes the model in terms of gradients and model parameters at two scales of time steps and tasks, effectively improving model performance.
The specific application of large models in Memory-Understanding based methods is to understand the representation of historical memory and to understand the environment and tasks based on its extensive world knowledge. NaviLLM [179] integrates the historical observation sequence into the embedding space through the visual encoder, inputs the multi-modal information of the fusion encoding into the large language model and fine-tunes it, reaching the state-of-the-art on multiple benchmarks. NaVid [184] makes improvements in the encoding of historical information, achieves different degrees of information retention on historical observations and current observations through different degrees of pooling. DiscussNav [181] assigns large model experts with different abilities to different roles, drives the large models to discuss before navigation actions to complete navigation decisions, and achieves excellent performance in zero-shot VLN.
@@
Some reinforcement learning methods are also applied to predict and explore future states. LookBY [185] employs reinforcement prediction to enable the prediction module to imitate the world and forecast future states and rewards. This allows the agent to directly map “current observations” and “predictions of future observations” to actions, achieving state-of-the-art performance at the time. The rich world knowledge and zero-shot performance of large models provide many possibilities for Future-Prediction based methods. MiC [188] requires the large language model to directly predict the target and its possible location from the instructions and provides navigation instructions through the description of scene perception. This method requires LLMs to fully exert its ‘imagination’ and build an imagined scene through prompts.
@@
IV-D Non-Visual Perception
Tactile sensors provides agents with detailed information about object properties such as texture, hardness and temperature. It allows robots to fully grasp the high-precision tasks in hand, making it vital for agents in the physical world. Tactile perception undoubtedly enhances human-computer interaction and holds great promise [193, 194, 195].
For tactile perception tasks, the agent needs to gather tactile information from the physical world and then perform complex tasks. In this section, as shown in Fig. 10, we first introduce the existing types of tactile sensors and their datasets, and then discuss three main tasks in tactile perception: estimation, recognition, and manipulation.
@@ IV-D1 Sensor Design
The principle of human tactile is that the skin changes shape when touched, and its abundant nerve cells send electrical signals [196],which also serves as the basis for designing tactile sensors. Tactile sensor design methods can be divided into three categories: non-vision-based, vision-based, and multi-modal. Non-vision-based tactile sensors, primarily using electrical and mechanical principles, chiefly register fundamental, low-dimensional sensory outputs such as force, pressure, vibration, and temperature[197, 198, 199, 200, 201, 202]. One of the notable representatives is BioTac [203] and its simulator[204]. Vision-based tactile sensors are based on optical principles. Using images of the gel’s deformation as tactile information, vision-based tactile sensors such as GelSight[205], Gelslim[206], DIGIT[207], 9DTact[208], TacTip[209], GelTip[210] and AllSight[211] have been used for numerous applications. Simulations like TACTO[212] and Taxim[213] are also popular. Recent work has focused on cost reduction[208] and integration into robotic hands[207, 214, 215]. Multi-modal tactile sensors, inspired by human skin, combine multi-modal information like pressure, proximity, acceleration, and temperature, using flexible materials and modular design.
@@ IV-D2 Datasets
The datasets of non-vision sensors contain electrode values 3D net force vectors and contact location. Therefore, the objects in the datasets are usually force samples and grasping samples. Its tasks are mainly the estimation of force types, force value and grasping details. The datasets mainly collected by BioTac series[203]. As for vision-based sensors, with their high-resolution images of deformation gel, except estimating the force information and sliding, they focus more on texture classification and 3D reconstruction. The objects in the datasets are usually household object, wildlife environments, different materials and grasping items.At the same time, as the image information can be easily aligned and bound with the other modalities (images, language, audio, etc)[16, 216], the perception of touch in embodied agent mainly revolves around visual-based sensors. The datasets revolves around the Geisight sensors, DIGIT sensor and their simulators[205, 207, 212, 208]. We introduce ten commonly used tactile datasets, which are summarized in Table VIII.
TABLE VIII: Comparison of different vision-based tactile datasets.
@@ IV-D3 Methods
Tactile perception has numerous applications and can be categorized into three types: estimation, precise robotic manipulation, and multi-modal recognition tasks.
Estimation
Early works in estimation predominantly focused on basic algorithm for shape, force and slip measurement [225, 226, 208]. Researchers simply used a threshold or applied Convolutional Neural Networks (CNN) to settle these task, based on the colors of tactile images and the change of markers distributions at different frames. The emphasis of the estimation work is mainly on the second stage, the generation of tactile image and the reconstruction of the object. The generation of tactile image [227, 228, 229, 230] aims at generating tactile image from vision data. At first it applied deep learning model that takes RGB-D images as input and outputted tactile images [227, 228]. Recently, as the image generation developed fast, Higuera et al. [229] and Yang et al. [230] applied diffusion model on tactile generation and it performed well. The reconstruction of the object can be divided into 2D reconstruction [231, 232] and 3D reconstruction [233, 234, 208, 235, 236, 237, 238, 239, 240, 241, 242, 243, 244, 245, 246, 247]. 2D reconstruction mainly focuses on the shape and segmentation of the object while 3D focuses on the surface and pose, and then even the full scene perception. The tasks first adopted mathematics methods, auto-encoder methods and neural network methods to fuse visual (sometime point clouds) and tactile features together. Lately, researchers like Comi et al. [241] and Dou et al. [242] implemented new methods based on Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS) into tactile reconstruction work.
Robotic Manipulation
In these task, bridging the sim-to-real gap is more than important. Reinforcement learning and GAN-based methods have been proposed to address variations in accurate, on-time robotic manipulation tasks.
@@
GAN-based method. ACTNet[252] proposed an unsupervised adversarial domain adaptation method to narrow the domain gap for pixel-level tactile perception tasks. An adaptively correlative attention mechanism was introduced to improve the generator, which is capable of leveraging global information and focusing on salient regions. However, pixel-level domain adaptation lead to error accumulation, degrade performance, and increased structural complexity and training costs. In comparison, STR-Net[253] proposed a feature-level unsupervised framework for tactile images, narrowing the domain gap for feature-level tactile perception tasks. Moreover, some methods focuse on sim-to-real. For example, the Tactile Gym 2.0 [254]. However, due to its complexity and high cost, it is challenging for practical application.
Recognition
Tactile representations learning focuses on material classification and multi-modal understanding, which can be divided into two categories: Traditional Methods and LLMs&VLMs Methods.
@@ IV-D4 Difficulties
a) Disadvantages of sensors with different principles: traditional sensors provide simple, limited, and low-dimensional data, posing challenges for multi-modal learning. Vision-based sensors and electronic skins, although offering high accuracy, are cost-prohibitive. Moreover, vision-based sensors cannot provide temperature information. b) Difficulties in data acquisition: tactile datasets are rare and heterogeneous, lacking the standardized, extensive repositories found in fields like vision. Data collection is particularly challenging, especially when gathering both tactile and visual information simultaneously, despite some efforts in developing simplified collection devices. c) Difficulties in inconsistent standards: there are various sensors on the market with inconsistent standards and principles. Even with similar imaging patterns, the design and calibration of vision-based tactile sensors result in a significant domain gap. Unlike standardized formats for visual and auditory data, tactile data formats vary widely among sensor manufacturers, making large-scale learning from heterogeneous sensors difficult and limiting the usefulness of publicly available tactile datasets.available tactile datasets.
@@ V-A Embodied Question Answering
For EQA task, the agent needs to explore the environment from a first-person perspective to gather information necessary to answer the given questions. An agent with autonomous exploration and decision-making capabilities must not only consider which actions to take to explore the environment but also determine when to stop exploring to answer questions. Existing works focus on different types of questions, some of which are shown in Fig. 11. In this section, we first introduce the existing datasets and then discuss related methods.
@@ V-A1 Datasets
SQA3D[270] simplified protocol (QA only) while still preserving the function of benchmarking embodied scene understanding, enabling more complex, knowledge-intensive questions and a much larger scale of data collection. Specifically, SQA3D offers a dataset with about 6,800 unique situations, 20,400 descriptions, and 33,400 diverse reasoning questions for these situations based on ScanNet[278] scenes.
@@ V-A2 Methods
The EQA task was first proposed by Das et al.[265]. In their work, the agent consists of four main modules: vision, language, navigation, and answering. These modules are primarily constructed using traditional neural building blocks: Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs). These modules undergo training in two phases. Initially, the navigation and response modules are trained independently on automatically generated expert navigation demonstrations using imitation or supervised learning. Subsequently, in the second phase, the navigation architecture is fine-tuned using policy gradients. Some subsequent works[280, 281] retain modules like the question answering module proposed by Das et al.[265] and improve the model. Additionally, Wu et al.[281] proposed integrating the navigation and QA modules into a unified SGD training pipeline for joint training, thereby avoiding employing deep reinforcement learning to simultaneously train the separately trained navigation and question answering modules.